OpenStudy (anonymous):

Arrange 6x2 – 2x4 + 4 + 3x in order by decreasing powers of x. A. 6x2 + 4 + 3x – 2x4 B. –2x4 + 6x2 + 3x + 4 C. –2x4 + 3x + 6x2 + 4 D. 4 + 3x + 6x2 – 2x4

OpenStudy (anonymous):

B @freckles @mathway @mathstudent55

OpenStudy (mathstudent55):

\(6x^2 – 2x^4 + 4 + 3x =\) Keep in mind that x is the same as \(x^1\) Also, since \(x^0 = 1\), \(4 = 4x^0\) So our expression is the same as \(=6x^2 – 2x^4 + 4x^0 + 3x^1 \) \(=-2x^4 + 6x^2 + 3x + 4\) You are correct.

OpenStudy (anonymous):

thank you.